Two vectors of magnitudes 3,4 acting at an angle 60.find the resultant

Grade:11

4 Answers

sreehari
6 years ago
r=√a^2+b^2+2abcos(θ)
The resultant r can be found out by simply substituting in the above equation,
Here r=√37
DIWAKAR PANDEY
6 years ago
By using r=√a^2+b^2+2abcos(theta) R=√9+16+2×3×4×cos60 R=√25+12 R=√37Ans=.......................R=√37
Aditya kashyap
6 years ago
Using ||gm law of vector addition ... R =√( a^2+b^2+2abcos@)...Here a =3 , b = 4, @= 60°.... R =√ ( 3^2 +4^2 +2*3*4cos60°)... R=√(9+ 16+24*1/2)......R=√37 .....
Kusumanchi Vinay
6 years ago
R = ROOT OF A^2+B^2+2ABcostheta …. A=3 and B=4....and theta=60....R=root(9+16+24*1/2) .....So, R= root37